Tom reads at an average rate of 30 pages per hour, while Jan reads at an average rate of 40 pages per hour. If Tom starts reading a novel at 4:30 and Jan begins reading an identical copy of the same book at 5:20, at what time will they be reading the same page?
Answer by lwsshak3(11628)

Tom reads at an average rate of 30 pages per hour, while Jan reads at an average rate of 40 pages per hour. If Tom starts reading a novel at 4:30 and Jan begins reading an identical copy of the same book at 5:20, at what time will they be reading the same page?
**
30 pages/hr=30/60 pages/min=1/2 pages/min
40 pages/hr=40/60 pages/min=2/3 pages/min
5:20-4:30=50 min
..
let x=minutes Tom was reading novel
x-50=minutes Jan was reading novel
..
pages tom read=x*1/2=x/2
pages Jan read=2/3*(x-50)
x/2=2/3*(x-50)=(2x-100)/3
LCD=6
3x=2(2x-100)=4x-200
x=200 min=3hrs+20 min
ans:
Time Tom and Jan will be reading from the same page=4:30+3hrs+10 min=7:40